i need the graph for y<2x+4 and -3x-2y>=6
Answer by lwsshak3(11628)

i need the graph for y<2x+4 and -3x-2y>=6
**
I don't have the means to graph for you, but I can show you what must be done.
..
y<2x+4
This equation says y is less than all the values below the straight line, 2x+4.
To graph, draw a dashed line which has a slope of 2 and y-intercept=4, or plot points.
The dashed line means y-values on the line are not included in the inequality. To finish the graph, shade the area below the dashed line which covers all the y-values which satisfy the inequality
..
-3x-2y≥6
-2y≥3x+6
divide by -2 and reverse inequality sign
y≤-3x/2-3
This equation says y is less than all the values below and on the straight line, -3x/2-3
To graph, draw a solid line which has a slope of -3/2 and y-intercept=-6, or plot points.
The solid line means y-values on the line are included in the inequality. To finish the graph, shade the area below the solid line which covers all the y-values which satisfy the inequality
..
Finding point of intersection:
2x+4=-3x/2-3
4x+8=-3x-6
7x=-14
x=-2
y=2x+4=-4+4=0
point of intersection (-2,0)
If you draw your graphs correctly, the lines will cross at these coordinates and you will have overlapping of areas starting from this point.